1. Consider the ideal steam cycle shown below that incorporates both reheat and a closed feedwater heater (CFH). a. CompleteState T°F) P (psia) h (BTU/ lbm) s (BTU/lbm R) × (Quality) Relative Mass Flow (m/m) 2s 3 4 5 6 7s 8s 9 10
Answer #1

- → har = hochs TS=T4 (Ps 6310f TPA lalo = horby = cp (T6-7 Ty z To Pro 1000 60 TY - T7= 1 P6 673°F The total work done in tuPin 24.15 Thermal efficiency – Wtotal = 181.9. n = 0.764 WT Din = 400, oro z 523. 56 MW - Qin 0.764 in z 523.56410 J 2 4962 3